OpenStudy (anonymous):

A line passes through the point(3,-8) and has a slope of -6 . Write an equation for this line.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

A line passes through the point(3,-8) and has a slope of -6 . Write an equation for this line.

OpenStudy (richyw):

\[y=mx+b\]the slope is \(-6\), so \(m=6\)now you know that it passes through the point where \(x=3\) and \(y=-8\). So you just need to solve for \(b\).\[-8=-6(3)+b\]\[-8=-18+b\]\[b=10\]

OpenStudy (richyw):